# Maneuvering measurement-coherence into measurement-entanglement

### v2

@@ Abstract
Quantum dynamics governs the transformation of static quantum resources, such as coherence and entanglement, in both quantum states and measurements. Prior studies have established that a quantum channel’s state-cohering power can be converted into state-entangling power without additional coherence. Here, we complete this coherence-to-entanglement paradigm by demonstrating that a channel’s measurement-cohering power can likewise be converted into measurement-entangling power. This result reinforces, on the dynamical level, the intuition that entanglement emerges as a manifestation of coherence. To formalize this picture, we develop resource theories for measurement-cohering and measurement-entangling powers and characterize the structure of incoherent measurements to analyze measurement-coherence generation. Furthermore, we show that the state-cohering power of a quantum channel is equivalent to the measurement-cohering power of its adjoint map, and a corresponding equivalence also exists between state-entangling power and measurement-entangling power.
Introduction.—Quantum coherence, deeply rooted in the linear structure of quantum theory [1], characterizes quantum superposition and underlies iconic phenomena such as Schrödinger’s cat and the double-slit interference. Its role extends broadly across diverse quantum technologies, including quantum sensing [2, 3], quantum computation [4, 5, 6], quantum correlations [7], and quantum thermodynamics [8, 9]. Quantum coherence has thus been rigorously formalized within quantum resource theory [10, 11, 12, 13, 14], paralleling the framework initially developed for entanglement. Quantum coherence is also a prerequisite for quantum entanglement, which can be intuitively understood as coherence distributed among multiple quantum subsystems. This fundamental connection has inspired extensive studies exploring the conversion of state-coherence into state-entanglement under incoherent operations within continuous-variable optical systems [15, 16, 17, 18] as well as finite-dimensional systems [19, 20, 21, 22].
This coherence-entanglement relationship extends beyond quantum states to quantum measurements. Quantum measurements themselves hold independent significance as carriers of quantum resources [23, 24, 25, 26, 27, 28, 29, 30]. Quantum coherence in measurements—measurement-coherence—is indispensable for comprehensively probing quantum states with coherence [27, 31, 32]. Similarly, quantum entanglement in measurements—measurement-entanglement—offers advantages in quantum information tasks such as quantum state discrimination and quantum system learning [28, 33, 34]. Recently, measurement-coherence has also been shown capable of converting into measurement-entanglement [35].
@@
Figure 1: Quantum dynamics can modify static quantum resources either in quantum states or in quantum measurements.
Quantum dynamics, equally fundamental in quantum theory alongside states and measurements, plays a more intricate role concerning quantum resources, as depicted in Fig. 1. Quantum dynamics can create [36, 37, 38], preserve [39], or destroy static quantum resources [40]. Beyond merely altering resources, dynamics themselves constitute independent quantum resources capable of simulating other processes that lack intrinsic quantum resources [41, 42]. While the interplay between quantum dynamics and static quantum resources significantly enriches resource theories, notably, it has recently been demonstrated that certain aspects of quantum dynamics related to static quantum resources parallel the relationship between static coherence and entanglement. Specifically, it has been shown that the state-cohering power of quantum dynamics can be converted into state-entangling power [43].
In this Letter, we complete this foundational coherence-entanglement relationship at the dynamical level by introducing and analyzing a resource theory for quantum channels based on their ability to generate measurement-resources, specifically measurement-cohering power and measurement-entangling power. We establish a fundamental relation showing that the measurement-cohering power of quantum dynamics can be converted into measurement-entangling power. To achieve these results, we demonstrate that all incoherent measurements reduce to canonical incoherent basis measurements. Furthermore, we reveal an equivalence between the state-cohering power of a unital quantum channel and the measurement-cohering power of its adjoint channel, accompanied by a corresponding equivalence for entanglement. Our findings thus provide a unified and consistent picture of how quantum dynamics’ capacity to generate static coherence can be converted into the capacity to generate static entanglement, static resources in the form of either state-resources or measurement-resources, thereby significantly broadening the conceptual scope of quantum resource theories.
@@
Measurement-coherence and measurement-entanglement.—In the resource theory of quantum coherence [14], quantum coherence of an operator is quantified with respect to a fixed orthonormal basis, which is referred to as the incoherent basis. The choice of an incoherent basis is contingent upon the experimental or physical constraints imposed on the system under consideration. We set an incoherent basis of a system $A$ and denote it as $\mathcal{I}_{A}=\{\lvert i\rangle_{A}:i=0,\dots,d-1\}$, with respect to which quantum coherence is considered. An operator $O_{A}$ that is diagonal in the incoherent basis is referred to as incoherent. Any incoherent operator is a fixed point under the dephasing channel $\Delta_{A}(O_{A})=\sum_{i=0}^{d-1}\langle i\rvert O_{A}\lvert i\rangle_{A}\left|i\right\rangle\!\left\langle i\right|_{A}=O_{A}$. Similarly, a quantum state $\rho_{A}$ is incoherent if $\Delta_{A}(\rho_{A})=\rho_{A}$. In this Letter, we consider constituent systems of dimension $d$ and assume that a quantum measurement for each system has $n$ outcomes, which can be described as a positive operator valued measure (POVM) given by $\mathcal{M}_{A}=\{M_{x}:\sum_{x=0}^{n-1}M_{x}=I_{A},M_{x}\geq 0\;\forall x=0,\ldots,n-1\}$, where $I_{A}$ is the identity operator. A quantum measurement $\mathcal{M}_{A}$ can be equivalently described by a quantum channel with classical outputs as
@@
where $\rho_{A}$ is a quantum state and $A^{\prime}$ denotes the output system [44]. The output of a quantum measurement is regarded as being classical implying that $\{\left|x\right\rangle\!\left\langle x\right|_{A^{\prime}}:x=0,\dots,n-1\}$ is an incoherent basis for the system $A^{\prime}$. A quantum measurement $\mathcal{M}_{A}$ is called an incoherent measurement if its POVM elements are incoherent, that is, $\Delta_{A}(M_{x})=M_{x}$ for all $x=0,\dots,n-1$. If a quantum measurement $\mathcal{M}_{A}$ is incoherent, the probability of an outcome is solely dependent on the incoherent information of input states, disregarding the coherent components [27, 45], as demonstrated by the following equation:
$p(x)=\operatorname{Tr}_{A}M_{x}\rho_{A}=\operatorname{Tr}M_{x}\Delta_{A}(\rho_{A}).$ (2)
We denote the set of incoherent measurements as $\mathbf{IM}$. Measurement-coherence of a quantum measurement $\mathcal{M}_{A}$ is quantified by the relative entropy of measurement-coherence $C_{m}(\cdot)$, as defined in [35]:
@@
Figure 2: A quantum measurement $\mathcal{M}_{A}$ transforms to another quantum measurement ~$\widetilde{\mathcal{M}}_{A}$ under a pre-processing channel $\mathcal{E}_{A}$ and a classical post-processing channel $\mathcal{S}_{A^{\prime}}$. The pre-processing channel can generate measurement-resource.
A quantum measurement $\mathcal{M}_{A}$ undergoes a transformation into another quantum measurement ~$\widetilde{\mathcal{M}}_{A}$ when a pre-processing channel and a classical post-processing channel are added, as illustrated in Fig. 2 [23]. This can be described as follows:
~$\widetilde{\mathcal{M}}_{A}=\mathcal{S}_{A^{\prime}}\circ\mathcal{M}_{A}\circ\mathcal{E}_{A},$ (6)
@@
We defer proofs of the above and subsequent results to the Supplemental Material. The set of quantum channels that do not generate any measurement-coherence is known to be detection-incoherent operations ($\mathbf{DIO}$) [27]. These serve as free operations for measurement-coherence. A quantum channel $\mathcal{E}_{A}$ is in $\mathbf{DIO}$ if and only if it satisfies the equation $\Delta_{A}\circ\mathcal{E}_{A}=\Delta_{A}\circ\mathcal{E}_{A}\circ\Delta_{A}$. This characterization is analogous to the characterization of a quantum channel in the maximal incoherent operations ($\mathbf{MIO}$), which is the largest set of quantum channels that send incoherent states to incoherent states.
In the resource theory of measurement-entanglement, we take the closure of the set of LOCC measurements ($\overline{\mathbf{LOCCM}}$) as our free resources, that can be implemented as LOCC operations with local measurements [44]. The closure of the set of LOCC channels ($\overline{\mathbf{LOCC}}$) corresponds to the free operations that preserve $\overline{\mathbf{LOCCM}}$. Measurement-entanglement of a bipartite quantum measurement $\mathcal{M}_{AB}$ is quantified by the relative entropy of measurement-entanglement of $\mathcal{M}_{AB}$ defined as follows  [^1] We choose $\overline{\mathbf{LOCCM}}$ as our free resource, which possesses a more direct operational meaning than $\mathbf{SepM}$, the free resource used in [35].:
@@
We remark that, despite the similarities between the resource theories of quantum coherence and entanglement, there is no entanglement-destroying channel that destroys entanglement while holding every separable state as its fixed point [47]. This is in contrast to the case of quantum coherence, where the dephasing channel destroys static quantum coherence while holding every incoherent state as a fixed point of the channel. As a result of this limitation, there is no straightforward operational interpretation or concise formula for $E_{m}$ as there is for the relative entropy of measurement-coherence reduced in Eq. (5).
Measurement-resource powers of quantum dynamics.—Now we focus on the capacity of quantum dynamics to generate measurement-resources. Quantum channels are capable of generating measurement-coherence and measurement-entanglement by acting as a pre-processing channel prior to a quantum measurement. In contrast, classical post-processing after a quantum measurement is unable to generate any measurement-resources.
As previously introduced, $\mathbf{DIO}$ is the largest set of quantum channels under which incoherent measurements remain incoherent [35]. In light of this, we define the measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ as follows:
where the ancillary system $\widetilde{A}$ can be arbitrary. In fact, we find that the measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ is equal to the amount of the measurement-coherence generated by the quantum channel $\mathcal{E}_{A}$ from the incoherent basis measurement $\mathcal{I}_{A}$:
@@
Similarly, the free resource in measurement-entangling power corresponds to $\overline{\mathbf{LOCC}}$ channels. We define the measurement-entangling power of a quantum channel $\mathcal{E}_{AB}$ as follows:
@@
where the ancillary systems $\widetilde{A}$ and $\widetilde{B}$ can be arbitrary. As previously noted, there is no resource-destroying dynamics that fixes any separable state in entanglement theory. Thus, the aforementioned simplification as in Eq. (10) is not applicable.
@@
Conversion of measurement-cohering power to measurement-entangling power.—Having established the concepts of measurement-cohering and measurement-entangling powers of quantum channels, we now present our main result: quantum coherence can be converted to entanglement at the level of dynamics generating measurement resources as illustrated in Fig. 3(a). First, we demonstrate that the measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ serves as an upper bound on the measurement-entangling power of a composite channel constructed from the quantum channel $\mathcal{E}_{A}$, without any additional measurement-cohering power:
@@
Let $\mathcal{E}_{A}$ be a quantum channel. Then we have that
@@
for a $\mathbf{DIO}$ channel $\mathcal{K}_{AB}$ and a unital $\mathbf{DIO}$ channel $\mathcal{L}_{AB}$.
Given that the quantum channels surrounding the quantum channel $\mathcal{E}_{A}$ are $\mathbf{DIO}$s, they do not contribute measurement-cohering power (See Fig. 3(a)). The dephasing channels in the post-processing serve to render the subsequent measurement incoherent and align the incoherent basis, thereby enabling the foregoing channels to convert the measurement-cohering power of the quantum channel $\mathcal{E}_{A}$ into the measurement-entangling power in that incoherent basis. In contrast, when converting state-cohering power to state-entangling power, dephasing channels serve as a preprocessing step to ensure the input states are incoherent, as depicted in Fig. 3(b) [43]. Consequently, Theorem 3 implies that measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ is the sole source of measurement-entangling power of the composite quantum channel. Indeed, there exists a configuration that allows for complete conversion of measurement-cohering power to measurement-entangling power although we restrict the pre-processing channel $\mathcal{L}_{AB}$ to be unital, as the following result shows (see Table 1):
@@
For a quantum channel $\mathcal{E}_{A}$, its measurement-cohering power can be fully converted to measurement-entangling power under a pre-processing by the adjoint channel of the generalized CNOT channel $\mathcal{U}_{\mathrm{CNOT}}^{{\dagger}}$ and a post-processing dephasing channel:
@@
Figure 3: (a) Measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ can convert to measurement-entangling power of a quantum channel ~$\widetilde{\mathcal{E}}_{AB}$ under a $\mathbf{DIO}$ channel $\mathcal{K}_{AB}$, a unital $\mathbf{DIO}$ channel $\mathcal{L}_{AB}$, and the dephasing channel $\Delta_{AB}$. (b) State-cohering power of a quantum channel $\mathcal{E}_{A}$ can convert to state-entangling power of a quantum channel ~$\widetilde{\mathcal{E}}_{AB}$ under $\mathbf{MIO}$ channels $\mathcal{K}_{AB}$ and $\mathcal{L}_{AB}$, and the dephasing channel $\Delta_{AB}$ [43].
This result completes the picture of quantum coherence convertible to quantum entanglement at the dynamic level, specifically by addressing the previously unexplored role of quantum channels on the half of the static resources: measurement-resources. Measurement resources constitute half of the static resources in quantum information—complementing the widely studied state resources—yet have received comparatively less attention. This critical balance between measurement and state resources is fundamental, as state-coherence itself is inherently incomplete without coherent measurement to observe its effects [27, 32]. Similarly, measurement-entanglement is crucial in achieving exponential quantum advantage in quantum learning [33, 49, 50, 34].
The convertibility of measurement-cohering power to measurement-entangling power implies that a measurement-entangling power monotone induces a measurement-cohering power monotone, as demonstrated by the following result:
@@
State-resource power versus measurement-resource power.—A quantum channel $\mathcal{E}_{A}$ can generate static quantum resources in both quantum states and measurements. The state-cohering power of $\mathcal{E}_{A}$, quantifying the maximum coherence the channel can induce in incoherent states, is given by
$C_{g}(\mathcal{E}_{A})\coloneqq\max_{\sigma_{A\widetilde{A}}\in\mathbf{I}}C_{R}(\mathcal{E}_{A}(\sigma_{A\widetilde{A}})),$ (16)

### v3

@@ Abstract
Quantum dynamics governs the transformation of static quantum resources, such as coherence and entanglement, in both quantum states and measurements. Prior studies have established that a quantum channel’s state-cohering power can be converted into the state-entangling power without additional coherence. Here, we complete this coherence-to-entanglement paradigm by demonstrating that a channel’s measurement-cohering power can likewise be converted into the measurement-entangling power. This result reinforces, on the dynamical level, the intuition that entanglement emerges as a manifestation of coherence. To formalize this picture, we develop resource theories for measurement-cohering and measurement-entangling powers and characterize the structure of incoherent measurements to analyze measurement-coherence generation. Furthermore, we show that the state-cohering power of a quantum channel is equivalent to the measurement-cohering power of its adjoint map, and a corresponding equivalence also exists between the state-entangling power and the measurement-entangling power.
1 Introduction
Quantum coherence is a fundamental manifestation of the superposition principle and lies at the heart of many quintessential quantum phenomena such as double-slit interference and Schrödinger’s cat. In recent decades, coherence has been formalized within the framework of quantum resource theories [1, 2, 3, 4, 5], leading to a unified understanding of its operational roles in quantum sensing [6, 7], quantum computation [8, 9, 10], and thermodynamics [11, 12]. In parallel, entanglement—another central quantum resource—has been extensively studied, and its deep connection to coherence has been clarified: coherence distributed across different subsystems naturally gives rise to entanglement. This connection has been explored both conceptually and operationally, yielding explicit protocols that convert state-coherence into state-entanglement under appropriate restrictions on quantum operations [13, 14, 15, 16, 17, 18, 19, 20].
While these developments have primarily focused on states, quantum measurements themselves also constitute carriers of non-classical resources [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. A quantum measurement can possess coherence with respect to a fixed basis, enabling the extraction of phase-sensitive information that cannot be accessed by incoherent measurements [27, 33, 34]. Likewise, measurements can exhibit entanglement, allowing joint measurement strategies that outperform any LOCC implementable scheme [29, 35, 36]. Despite their importance, measurement-based resources remain less explored than their state-based counterparts. Recent work has begun to fill this gap by establishing resource theories for measurement-coherence and measurement-entanglement and by demonstrating that measurement-coherence can be converted into measurement-entanglement in static scenarios [37].
@@ 1 Introduction
Figure 1: Quantum dynamics can modify static quantum resources in both quantum states and quantum measurements.
A complementary and equally fundamental aspect of quantum theory is quantum dynamics. Quantum channels do not merely transform states and measurements; they also serve as dynamical resources capable of generating [38, 39, 40], destroying [41, 42], or preserving static quantum resources [43] as depicted in Fig. 1. In addition, quantum channels themselves may function as dynamical resources, thereby facilitating the simulation of processes that do not possess intrinsic quantum resources [44, 45, 46, 47, 48]. Viewed from this dynamical perspective, coherence and entanglement appear not only in static objects such as states or POVMs but also in the transformations applied to them. A recent result identified a relation between such dynamicals resources: the state-cohering power of a quantum channel can be fully converted into its state-entangling power, revealing a dynamical analogue of the coherence–entanglement correspondence known for states [49].
However, the dynamical landscape for measurement resources has remained largely unexplored. In particular, it is not known whether measurement-cohering power—the ability of a channel to generate measurement-coherence through pre-processing—can be operationally related to measurement-entangling power, or whether these two quantities obey a dynamical correspondence analogous to the one established for states. Establishing such a relationship is essential for completing the resource-theoretic picture linking coherence and entanglement across both static and dynamical settings.
In this work, we develop resource theories for quantum channels that quantify their ability to generate measurement-coherence and measurement-entanglement, and we show that these two forms of dynamical resources are fundamentally connected. We first demonstrate that the measurement-cohering power of any quantum channel can, without introducing any additional coherence, be fully converted into measurement-entangling power by means of a fixed construction involving only free operations and the final dephasing step. This establishes a genuine dynamical analogue of coherence-to-entanglement conversion on the measurement side.
A useful ingredient in our analysis is a structural characterization of the incoherent measurements: we show that every incoherent POVM can be reduced to a standard incoherent basis measurement followed solely by classical post-processing. This observation allows us to treat incoherent measurements as canonical objects and provides the foundation for analyzing how quantum channels generate measurement-resources.
Finally, we uncover a duality between state-based and measurement-based dynamical resource powers. For unital channels, the state-cohering power of the channel turns out to coincide with the measurement-cohering power of its adjoint map, and an analogous equivalence emerges between state-entangling and measurement-entangling powers. Taken together, these results complete the conceptual framework connecting coherence and entanglement across both static and dynamical settings, while highlighting a deeper symmetry between state- and measurement-based quantum resources and the unifying role played by quantum dynamics.
2 Measurement-coherence and measurement-entanglement
@@
In the resource theory of quantum coherence [5], quantum coherence of an operator is quantified with respect to a fixed orthonormal basis, which is referred to as an incoherent basis. The choice of incoherent basis is contingent upon the experimental or physical constraints imposed on the system under consideration. We set an incoherent basis of a system $A$ and denote it as $\mathcal{I}_{A}=\{\lvert i\rangle_{A}:i=0,\dots,d-1\}$, with respect to which quantum coherence is considered. An operator $O_{A}$ that is diagonal in the incoherent basis is referred to as incoherent. Any incoherent operator is a fixed point under the dephasing channel $\Delta_{A}(O_{A})=\sum_{i=0}^{d-1}\langle i\rvert O_{A}\lvert i\rangle_{A}\left|i\right\rangle\!\left\langle i\right|_{A}=O_{A}$. Similarly, a quantum state $\rho_{A}$ is incoherent if $\Delta_{A}(\rho_{A})=\rho_{A}$. Here, we consider constituent systems of dimension $d$ and assume that a quantum measurement for each system has $n$ outcomes, which can be described as a positive-operator-valued measure (POVM) given by $\mathcal{M}_{A}=\{M_{x}:\sum_{x=0}^{n-1}M_{x}=I_{A},M_{x}\geq 0\;\forall x=0,\ldots,n-1\}$, where $I_{A}$ is the identity operator. A quantum measurement $\mathcal{M}_{A}$ can be equivalently described by a quantum channel with classical outputs, as
@@
where $\rho_{A}$ is a quantum state and $A^{\prime}$ denotes the output system [50]. The output of a quantum measurement is regarded as being classical implying that $\{\left|x\right\rangle\!\left\langle x\right|_{A^{\prime}}:x=0,\dots,n-1\}$ is an incoherent basis for the system $A^{\prime}$. A quantum measurement $\mathcal{M}_{A}$ is called an incoherent measurement if its POVM elements are incoherent, that is, $\Delta_{A}(M_{x})=M_{x}$ for all $x=0,\dots,n-1$. If a quantum measurement $\mathcal{M}_{A}$ is incoherent, the probability of an outcome is solely dependent on the incoherent information of input states, disregarding the coherent components [27, 51]. We denote the set of incoherent measurements as $\mathbf{IM}$. Measurement-coherence of a quantum measurement $\mathcal{M}_{A}$ is quantified by the relative entropy of measurement-coherence $C_{m}(\cdot)$, as defined in Ref. [37]:
@@
A quantum measurement $\mathcal{M}_{A}$ undergoes a transformation into another quantum measurement $\widetilde{\mathcal{M}}_{A}$ when a pre-processing channel and a classical post-processing channel are added [21, 37]. This can be described as follows:
$\widetilde{\mathcal{M}}_{A}=\mathcal{S}_{A^{\prime}}\circ\mathcal{M}_{A}\circ\mathcal{E}_{A},$ (5)
@@
We defer proofs of the above and subsequent results to Appendices. The set of quantum channels that do not generate any measurement-coherence is known to be detection-incoherent operations ($\mathbf{DIO}$) [27]. These serve as free operations for measurement-coherence. A quantum channel $\mathcal{E}_{A}$ is in $\mathbf{DIO}$ if and only if it satisfies the equation $\Delta_{A}\circ\mathcal{E}_{A}=\Delta_{A}\circ\mathcal{E}_{A}\circ\Delta_{A}$. This characterization is analogous to the characterization of a quantum channel in the maximally incoherent operations ($\mathbf{MIO}$), which is the largest set of quantum channels that send incoherent states to incoherent states.
In the resource theory of measurement-entanglement, we take the closure of the set of LOCC measurements ($\overline{\mathbf{LOCCM}}$) as our free resources, that can be implemented as LOCC operations with local measurements [50]. The closure of the set of LOCC channels ($\overline{\mathbf{LOCC}}$) corresponds to the free operations that preserve $\overline{\mathbf{LOCCM}}$. Measurement-entanglement of a bipartite quantum measurement $\mathcal{M}_{AB}$ is quantified by the relative entropy of measurement-entanglement of $\mathcal{M}_{AB}$ defined as follows [^1] We choose $\overline{\mathbf{LOCCM}}$ as our free resource, which possesses a more direct operational meaning than the set of separable measurements, the free resource used in Ref. [37].:
@@
We remark that, despite the similarities between the resource theories of quantum coherence and entanglement, there is no entanglement-destroying channel that destroys entanglement while holding every separable state as its fixed point [52]. This is in contrast to the case of quantum coherence, where the dephasing channel destroys static quantum coherence while holding every incoherent state as a fixed point of the channel.
3 Measurement-resource powers of quantum dynamics
Now we focus on the capacity of quantum dynamics to generate measurement-resources. Quantum channels are capable of generating measurement-coherence and measurement-entanglement by acting as a pre-processing channel prior to a quantum measurement. In contrast, classical post-processing after a quantum measurement is unable to generate any measurement-resources [37].
As previously introduced, $\mathbf{DIO}$ is the largest set of quantum channels under which incoherent measurements remain incoherent. In light of this, we define the measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ as follows:
$C(\mathcal{E}_{A})\coloneqq$ $\min_{\mathcal{F}_{A}\in\mathbf{DIO}}\max_{\mathcal{M}_{A\widetilde{A}}\in\mathbf{IM}}$
$D_{m}(\mathcal{M}_{A\widetilde{A}}\circ\mathcal{E}_{A}\|\mathcal{M}_{A\widetilde{A}}\circ\mathcal{F}_{A}),$ (8)
where the ancillary system $\widetilde{A}$ can be arbitrary. The definition in Eq. (3) may appear formal, but it has a direct physical interpretation. The optimization over all incoherent measurements is unnecessary; the maximizing measurement is always the canonical incoherent basis measurement $\mathcal{I}_{A}$. Intuitively, the measurement-cohering power captures how much the quantum channel $\mathcal{E}_{A}$ “twists” the classical axes of the standard incoherent measurement $\mathcal{I}_{A}$ into a superposition, thereby generating coherence in the measurement process. This leads to the following operational simplification:
@@ 3 Measurement-resource powers of quantum dynamics
Theorem 2 ensures that the generalized Fourier transform corresponds to a quantum channel that has the maximum measurement-cohering power as in the case for the state-cohering power [46] since the generalized Fourier transform turns the incoherent basis measurement into a measurement with maximum measurement-coherence.
Furthermore, Theorem 2 establishes a direct pathway for experimental verification of the measurement-cohering power. Since the measurement-cohering power of a channel $\mathcal{E}$ is equivalent to the measurement-coherence of the effective measurement $\mathcal{I}\circ\mathcal{E}$, it opens the possibility of quantifying it using well-established techniques for characterizing static coherence. For instance, quantitative witnesses for coherence [53, 54], although originally developed for quantum states, might be adapted to quantify measurement-coherence, leveraging the fact that POVM elements are positive operators akin to unnormalized quantum states. This bridges our dynamical framework with existing experimental protocols for static resources.
Similarly, we can quantify the capability of a channel to generate measurement-entanglement. By identifying the closure of LOCC channels ($\overline{\mathbf{LOCC}}$) as the set of free resources, we define the measurement-entangling power of a quantum channel $\mathcal{E}_{AB}$ as:
@@
where the ancillary systems $\widetilde{A}$ and $\widetilde{B}$ can be arbitrary. In the definition of the measurement-entangling power, the ancillary systems $\widetilde{A}$ and $\widetilde{B}$ cannot be omitted, in contrast to Theorem 2 for the measurement-cohering power, because they play essential operational roles. Without these ancillary systems, the channel that attains the maximum measurement-entangling power would be the generalized CNOT gate. However, once ancillary systems are allowed, the generalized SWAP channel achieves the maximum measurement-entangling power, as it converts locally implemented Bell measurements into measurements with maximum measurement-entanglement, paralleling the situation for state-entangling power [55].
@@
4 Conversion of measurement-cohering power to measurement-entangling power
Having established the concepts of measurement-cohering and measurement-entangling powers of quantum channels, we now present our main result: quantum coherence can be converted into entanglement at the level of dynamics generating measurement resources as illustrated in Fig. 2. First, we demonstrate that the measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ serves as an upper bound on the measurement-entangling power of a composite channel constructed from the quantum channel $\mathcal{E}_{A}$, without any additional measurement-cohering power:
@@ 4 Conversion of measurement-cohering power to measurement-entangling power
Let $\mathcal{E}_{A}$ be a quantum channel. Then
@@
for any $\mathbf{DIO}$ channel $\mathcal{K}_{AB}$ and any unital $\mathbf{DIO}$ channel $\mathcal{L}_{AB}$.
Figure 2: The measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ can be converted to the measurement-entangling power of a quantum channel $\widetilde{\mathcal{E}}_{AB}$ by composing it with a $\mathbf{DIO}$ channel $\mathcal{K}_{AB}$, a unital $\mathbf{DIO}$ channel $\mathcal{L}_{AB}$, and the dephasing channel $\Delta_{AB}$.
Given that the channels surrounding the quantum channel $\mathcal{E}_{A}$ are $\mathbf{DIO}$ channels, they do not contribute to the measurement-cohering power (see Fig. 2). The dephasing channels in the post-processing serve to render the subsequent measurement incoherent and align the incoherent basis, thereby enabling the foregoing channels to convert the measurement-cohering power of the quantum channel $\mathcal{E}_{A}$ into the measurement-entangling power in that incoherent basis. In contrast, when converting the state-cohering power to the state-entangling power, dephasing channels serve as a pre-processing step to ensure the input states are incoherent, as depicted in Table 1 [49]. Consequently, Theorem 3 implies that the measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ is the sole source of the measurement-entangling power of the composite quantum channel. Indeed, there exists a configuration that allows for complete conversion of the measurement-cohering power to the measurement-entangling power although we restrict the pre-processing channel $\mathcal{L}_{AB}$ to be unital, as the following result shows (see Table 1):
@@
For a quantum channel $\mathcal{E}_{A}$, its measurement-cohering power can be fully converted to the measurement-entangling power under a pre-processing by the adjoint channel of the generalized CNOT channel $\mathcal{U}_{\mathrm{CNOT}}^{{\dagger}}$ and a post-processing dephasing channel:
@@
State-resource Measurement-resource
Static [Uncaptioned image] [Uncaptioned image]
Dynamic [Uncaptioned image] [Uncaptioned image]
Table 1: Illustration of the optimal conversion of quantum coherence into entanglement at different levels of resource theory. The diagrams, arranged from the top left to the bottom left in a clockwise direction, depict: (i) the optimal conversion of quantum coherence of a quantum state $\rho_{A}$ into entanglement of a quantum state $\widetilde{\rho}_{AB}$ [17]; (ii) the optimal conversion of measurement-coherence of a quantum measurement $\mathcal{M}_{A}$ into measurement-entanglement of a quantum measurement $\widetilde{\mathcal{M}}_{AB}$ [37]; (iii) the optimal conversion of the measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ into the measurement-entangling power of a quantum channel $\widetilde{\mathcal{E}}_{AB}$ (this work); and (iv) the optimal conversion of the state-cohering power of a quantum channel $\mathcal{E}_{A}$ into the state-entangling power of a quantum channel $\widetilde{\mathcal{E}}_{AB}$ [49].
This result completes the picture of quantum coherence convertible to quantum entanglement at the dynamical level, specifically by addressing the previously unexplored role of quantum channels on the half of the static resources: measurement-resources. Measurement resources constitute half of the static resources in quantum information—complementing the widely studied state resources—yet have received comparatively less attention. This critical balance between measurement and state resources is fundamental, as state-coherence itself is inherently incomplete without coherent measurement to observe its effects [27, 34]. Similarly, measurement-entanglement is crucial in achieving exponential quantum advantage in quantum learning [35, 57, 58, 36].
The convertibility of the measurement-cohering power to the measurement-entangling power implies that a measurement-entangling power monotone induces a measurement-cohering power monotone, as demonstrated by the following result:
@@
5 State-resource power versus measurement-resource power
A quantum channel $\mathcal{E}_{A}$ can generate static quantum resources in both quantum states and measurements. The state-cohering power of $\mathcal{E}_{A}$, quantifying the maximum coherence the channel can induce in incoherent states, is given by
$),C_{g}(\mathcal{E}_{A})\coloneqq\max_{\sigma_{A\widetilde{A}}\in\mathbf{I}}C_{R}(\mathcal{E}_{A}(\sigma_{A\widetilde{A}}$)), (15)